Shay10:

The equation of a circle whose center is at (4, 0) and radius is length 2^(3) is (X-4)^2 +y^2=2^3 (X+4)^2+y^2=12 (X-4)^2+y^2=12

snowflake0531:

\[(x-h)^2 + (y-k)^2 = r^2\] where the center is (h,k) and the radius is r

Shay10:

Okay.

snowflake0531:

By the way, wdym "the radius is 2^3" also, are there 4 choices? not 3?

Shay10:

That's it

snowflake0531:

Well, based on what I wrote above, which choice do you think it is

Shay10:

The first one

snowflake0531:

I think that too

snowflake0531:

Although it's weird... it should be radius squared... which would make it 2^6, but welp the first one makes the most sense i suppose

snowflake0531:

WAIT is it actually \(2\sqrt{3}\) ???

Shay10:

I don't understand

Shay10:

Is this geometry or precalculus

snowflake0531:

Is it possible for you to give me a screenshot of the question?

snowflake0531:

@shay10 wrote:
Is this geometry or precalculus
circle stuff is geometry- and you're in the class- how do you not know what class you're taking-

Shay10:

Idk

snowflake0531:

well, can you give me a screenshot of the question